Adv. Math. Econ. 7, 33-46 (2005)

Law invariant convex risk measures

Marco Frittelli^{1} and Emanuela Rosazza Gianin^{2}

1 Dipartimento di Matematica per le Decisioni. Università degli Studi di Firenze,

Italy. Via C. Lombroso, 6/17. 50134 Firenze, Italy 2 Corresponding author. Dipartimento di Matematica e Statistica, Università di

Napoli Federico II, Italy. Via Cinzia, 80126 Napoli, Italy

Received: August 4, 2004 Revised: November 6, 2004 JEL classification: G11, G12, G13 Mathematical Subject Classification (2000): 60G42, 60G44 Abstract. As a generalization of a result by Kusuoka (2001), we provide the representation of law invariant convex risk measures. Very particular cases of law invariant coherent and convex risk measures are also studied. Key words: convex risk measures; coherent risk measures; law invariant risk measures.